Rita sprints 40m to the right in 5.5s.
What was her average velocity in m/s?
1
Expert's answer
2020-09-29T09:40:20-0400

The magnitude:


"v=\frac{s}{t}=\frac{40}{5.5}=7.3\frac{m}{s}"

The direction: rightwards.